?(Yield to? maturity)?The Saleemi? Corporation's ?$1 ,000 bonds pay 11 percent interest annually and have 15 years until maturity. You can purchase the bond for ?$935.

a. What is the yield to maturity on this? bond?

b. Should you purchase the bond if the yield to maturity on a? comparable-risk bond is 13 ?percent?

a. The yield to maturity on the Saleemi bonds is ____?%. ? (Round to two decimal? places.)

b. You should/should not purchase the bonds because your yield to maturity on the Saleemi bonds is greater/less than the one on a comparable risk bond. ? (Select from the? drop-down menus.)
Answer #1

a) Calculation of Yield To Maturity (YTM)

Yield To Maturity (YTM) can be calculated as

(interest per annum + average other cost per annum) / average fund employed

Here,

interest per annum = $1,000 * 11% = $110

average other cost per annum = (Redemption price-current market price)/life remaining to maturity

= (1000-935) / 15 = $4.33

average fund employed = (Redemption price+current market price)/2

= (1000+935)/2 = $967.50

Yield To Maturity (YTM) = (110+4.33) / 967.50

= 0.1182

= 11.82%

b) Since the yield to maturity on the Saleemi bonds is less than the one on a comparable risk bond,you should not purchase the bonds.